# Tony Lindeberg

### v1

@@ 4.2 Purely spatial receptive fields based on affine Gaussian smoothing
The motivation why we overparameterize the effective variability in the shapes of the receptive fields in this way is for reasons of geometric interpretation with regard to viewing variations for a visual observer that may observe the same local surface patch from different distances and viewing directions. If the affine matrix $A$ in a locally linearized perspective viewing model
@@
is normalized such that it corresponds to an a local orthonormal projection, then by the requirement of covariance under the group of geometric image transformations, a variability in depth $Z$ directly maps to a variability in the spatial scaling factor $S_{x}$, which in terms of the receptive fields will correspond to a variability in the spatial scale parameter $s$. Similarly, a variability in the affine transformation matrix will map to a variability in the spatial covariance matrix $\Sigma$.
By differentiating the affine Gaussian kernel (1) with respect to the image coordinates $x=(x_{1},x_{2})^{T}$ and the parameters $s$ and the elements in the 2-D positive symmetric definite covariance matrix
@@
we get the explicit expressions for these derivatives according to Figure 6.555The computation of these expressions, as well as all the other generalized Hermite polynomials in this paper, have been performed in Wolfram Mathematica. By comparing these expressions, we can then verify the following infinitesimal relationships with respect to variations in the scale parameter $s$ and the elements $\Sigma_{11}$, $\Sigma_{12}$ and $\Sigma_{22}$ of the spatial covariance matrix $\Sigma$:
@@
With regard to the differential relationships over the parameters $s$, $\Sigma_{11}$ and $\Sigma_{22}$, the corresponding partial differential equations are parabolic and can thereby only be evolved in the corresponding positive directions. Hence, the expressions for $\partial_{s}$, $\partial_{\Sigma_{11}}$ and $\partial_{\Sigma_{22}}$ correspond to the notions of infinitesimal generators for multi-parameter semi-groups. The differential relationship over the parameter $\Sigma_{12}$ does, on the other hand, correspond to a hyperbolic partial differential equation, and could therefore be evolved in both directions.
Figure 7: Generalized Hermite polynomials as arising from derivatives of the joint spatio-temporal Gaussian kernel $T(x,t;\;s,\tau,v)$ according to (95), based on smoothing with a spatially isotropic Gaussian kernel, with respect to the image coordinates $x=(x_{1},x_{2})^{T}$ and the time variable $t$ up to order 2, as well as with respect to the spatial scale parameter $s$, the temporal scale parameter $\tau$ and the elements $v_{1}$ and $v_{2}$ of the velocity vector $v$.
@@ 4.3 Spatio-temporal receptive fields based on isotropic spatial smoothing
The corresponding spatio-temporal scale-space representation of any 2+1-D video sequence $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ is then given by
@@
Differentiating the spatio-temporal kernel $T(x,t;\;s,\tau,v)$ according to (95) with respect the spatial image coordinates $x=(x_{1},x_{2})^{T}$ and the time variable $t$ up to order 2, as well as with respect to the spatial scale parameter $s$, the temporal scale parameter $\tau$ and the elements $v_{1}$ and $v_{2}$ of the velocity vector $v$, leads to the partial derivatives shown in Figure 7.
@@
Here, the differential relationships over the parameters $s$ and $\tau$, the corresponding partial differential equations are parabolic, and can therefore only be evolved in the corresponding positive directions. Hence, the expressions for $\partial_{s}$ and $\partial_{\tau}$ correspond to the notions of infinitesimal generators for multi-parameter semi-groups. The differential relationship over the parameters $v_{1}$ and $v_{2}$ do, on the other hand, correspond to a hyperbolic partial differential equations, and could therefore be evolved in both directions.
[…]
Figure 8: Generalized Hermite polynomials as arising from derivatives of the spatio-temporal Gaussian kernel $T(x,t;\;s,\Sigma,\tau,v)$ according to (119), based on smoothing with a spatially anisotropic affine Gaussian kernel, with respect to the image coordinates $x=(x_{1},x_{2})^{T}$ and the time variable $t$ up to order 2, as well as with respect to the spatial scale parameter $s$, the elements $\Sigma_{11}$, $\Sigma_{12}$ and $\Sigma_{22}$ of the spatial covariance matrix $\Sigma$, the temporal scale parameter $\tau$ and the elements $v_{1}$ and $v_{2}$ of the velocity vector $v$. (The explicit expressions for $\partial_{x_{1}t}T(x,t;\;s,\Sigma,\tau,v)$, $\partial_{x_{2}t}T(x,t;\;s,\Sigma,\tau,v)$ and $\partial_{tt}T(x,t;\;s,\Sigma,\tau,v)$ have been omitted here, because they do not fit within two lines each.)
@@ 4.4 Spatio-temporal receptive fields based on affine Gaussian smoothing
with the corresponding spatio-temporal scale-space representation of any 2+1-D video sequence $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$: according to
@@
Differentiating the above spatio-temporal kernel $T(x,t;\;s,\tau,v)$ with respect the spatial image coordinates $x=(x_{1},x_{2})^{T}$ and the time variable $t$ up to order 2, as well as with respect to the spatial scale parameter $s$, the elements $\Sigma_{11}$, $\Sigma_{12}$ and $\Sigma_{22}$ of the spatial covariance matrix $\Sigma$, the temporal scale parameter $\tau$ and elements $v_{1}$ and $v_{2}$ of the velocity vector $v$, then gives the partial derivatives shown in Figure 8.
@@
For this joint combination of the computational structures in the two partially simplified receptive field models in Sections 4.2 and 4.3, for the differential relationships over the parameters $s$, $\Sigma_{11}$, $\Sigma_{22}$ and $\tau$, it holds that the corresponding partial differential equations are parabolic and can therefore only be evolved in the corresponding positive directions. Hence, the expressions for $\partial_{s}$, $\partial_{\Sigma_{11}}$, $\partial_{\Sigma_{22}}$ and $\partial_{\tau}$ correspond to the notions of infinitesimal generators for multi-parameter semi-groups. The differential relationships over the parameters $\Sigma_{12}$, $v_{1}$ and $v_{2}$ do, on the other hand, correspond to hyperbolic partial differential equations, and could therefore be evolved in both directions.
@@ 5 Macroscopic cascade structures for the generalized Gaussian derivative model
To avoid explicit integration with of the previous infinitesimal relationships between receptive field responses for different values of the filter parameters, an alternative way to model such relationships is in terms of macroscopic relations. For simplicity, we again start by deriving such relations for the receptive field models that are purely based on smoothing with Gaussian kernels, thus initially excluding the non-causal spatio-temporal receptive field model based on convolution with the time-causal limit kernel (4) over the temporal domain.
@@ 5.1.1 Purely spatial receptive fields based on affine Gaussian smoothing
Applied to the purely spatial smoothing operation corresponding to convolution with the affine Gaussian kernel, it therefore follows that the scale-space representation $L(\cdot;\;s_{2},\Sigma_{2})$ at a coarser spatial scale $(s_{2},\Sigma_{2})$ can be computed from the scale-space representation $L(\cdot;\;s_{1},\Sigma_{1})$ at any finer spatial scale $(s_{1},\Sigma_{1})$ according to
@@
Due to the linearity of the corresponding operators, this cascade smoothing property extends to receptive field responses defined by applying spatial derivative operators ${\cal D}_{x}$ to the spatial scale-space representation of the form
@@ 5.1.2 Joint spatio-temporal receptive fields based on affine Gaussian smoothing
To express a corresponding cascade smoothing property over the joint spatio-temporal domain, let us make use of the fact that the composed effect of the smoothing with a a 2-D either isotropic Gaussian kernel or an anisotropic affine Gaussian kernel over the spatial domain in combination with smoothing with a 1-D non-causal temporal Gaussian kernel over the temporal domain can equivalently be obtained by smoothing with a joint spatio-temporal Gaussian kernel over the 2+1-D joint spatio-temporal domain. Thus, when using an anisotropic affine spatial Gaussian kernel, leading to a joint spatio-temporal kernel of the form (119)
@@
this operation can equivalently be described as a 2+1-D convolution with a joint spatio-temporal kernel $g_{2+1\text{-D}}\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}^{3}\times{\mathbb{S}}_{+}^{3}\rightarrow{\mathbb{R}}$ of the form
$g_{2+1-D}(p;\,m_{2+1\text{-D}},\Sigma_{2+1\text{-D}})=\\ =\frac{1}{(2\pi)^{3/2}\sqrt{\det{\Sigma_{2+1\text{-D}}}}}\,e^{-(p-m_{2+1\text{-D}})^{T}\Sigma_{2+1\text{-D}}^{-1}(p-m_{2+1\text{-D}})}$ (135)
@@
$\begin{split}\Sigma_{2+1\text{-D}}=\frac{\int_{p\in{\mathbb{R}}^{2+1}}p\,p^{T}\,T(p;\;s,\Sigma,\tau,v)\,\,dp}{\int_{p\in{\mathbb{R}}^{2+1}}T(p;\;s,\Sigma,\tau,v)\,dp}-m_{2+1\text{-D}}^{2},\end{split}$ (138)
@@
Thus, let us assume that we have two configurations of parameter settings that would correspond to a cascade smoothing property over the filter parameters for two joint spatio-temporal smoothing kernels of the form (170) with
  • the joint convolution kernel $T(\cdot,\cdot;\;s_{i},\Sigma_{i},\tau_{i},v_{i})$ in the lower layer with $v_{i}=(v_{i,1},v_{i,2})^{T}$ and the elements of the spatial covariance matrix $\Sigma_{i}$ being $\Sigma_{i,11}$, $\Sigma_{i,12}$ and $\Sigma_{i,22}$, and with
  • the joint convolution kernel $T(\cdot,\cdot;\;s_{j},\Sigma_{j},\tau_{j},v_{j})$ in the higher layer with $v_{j}=(v_{j,1},v_{j,2})^{T}$ and the elements of the spatial covariance matrix $\Sigma_{j}$ being $\Sigma_{j,11}$, $\Sigma_{j,12}$ and $\Sigma_{j,22}$.
@@
  • an incremental joint spatio-temporal convolution kernel $T(\cdot,\cdot;\;\Delta s,\Delta\Sigma,\Delta\tau,\Delta v)$ with $\Delta v=(\Delta v_{1},\Delta v_{2})^{T}$ and the elements of the spatial covariance matrix $\Delta\Sigma$ being $\Delta\Sigma_{11}$, $\Delta\Sigma_{12}$ and $\Delta\Sigma_{22}$,
@@
thus implying the explicit cascade smoothing property of the form
@@
Solving for the parameters $\Delta s$, $\Delta\Sigma_{11}$, $\Delta\Sigma_{12}$, $\Delta\Sigma_{22}$, $\Delta\tau$, $\Delta v_{1}$ and $\Delta v_{2}$ of the incremental spatio-temporal convolution kernel $T(\cdot,\cdot;\;\Delta s,\Delta\Sigma,\Delta\tau,\Delta v)$ then gives
@@ 5.1.3 Joint spatio-temporal receptive fields based on isotropic Gaussian smoothing
In relation to the results from the previous section, it is worth observing that if the spatial smoothing operation would instead only be performed based on an isotropic Gaussian kernel with the spatial covariance matrix $\Sigma$ equal to the unit matrix $\Sigma=I$, then the parameters $(s,\tau,v)$ in the spatio-temporal receptive fields would only span 4 out of the general 6 degrees of freedom in the joint spatio-temporal covariance matrix $\Sigma_{2+1\text{-D}}$ according to (139). Thereby, the configurations for which it would be possible to implement a corresponding cascade smoothing property
@@
would on the other hand be highly degenerate. Thus, beyond the previously mentioned closedness property under anisotropic spatial affine transformations, the use of an affine Gaussian kernel for spatial smoothing also offers clear advantages in terms of the ability for an idealized vision system to implement spatio-temporal cascade smoothing properties for the receptive fields, to decrease the amount of computational work in filter bank implementations of joint spatio-temporal receptive fields.

### v2

@@ 4.2 Purely spatial receptive fields based on affine Gaussian smoothing
The motivation why we overparameterize the effective variability in the shapes of the receptive fields in this way is for reasons of geometric interpretation with regard to viewing variations for a visual observer, that may observe the same local surface patch from different distances and viewing directions. If the affine matrix $A$ in a locally linearized perspective viewing model
@@
is normalized such that it corresponds to an a local orthonormal projection, then by the requirement of covariance under the group of geometric image transformations, a variability in depth $Z$ directly maps to a variability in the spatial scaling factor $S_{x}$, which in terms of the receptive fields will correspond to a variability in the spatial scale parameter $s$. Similarly, a variability in the affine transformation matrix $A$ will map to a variability in the spatial covariance matrix $\Sigma$.
By differentiating the affine Gaussian kernel (1) with respect to the image coordinates $x=(x_{1},x_{2})^{T}$ and the parameters $s$ and the elements $\Sigma_{i,j}$ in the 2-D positive symmetric definite covariance matrix
@@
we get the explicit expressions for these derivatives according to Figure 6.555The computations of these expressions, as well as all the other generalized Hermite polynomials in this paper, have been performed in Wolfram Mathematica. By comparing these expressions, we can then verify the following infinitesimal relationships with respect to variations in the scale parameter $s$ and the elements $\Sigma_{11}$, $\Sigma_{12}$ and $\Sigma_{22}$ of the spatial covariance matrix $\Sigma$:
@@
With regard to the differential relationships over the parameters $s$, $\Sigma_{11}$ and $\Sigma_{22}$, the corresponding partial differential equations are parabolic, and can thereby only be evolved in the corresponding positive directions. Hence, the expressions for $\partial_{s}$, $\partial_{\Sigma_{11}}$ and $\partial_{\Sigma_{22}}$ correspond to the notions of infinitesimal generators for multi-parameter semi-groups. The differential relationship over the parameter $\Sigma_{12}$ does, on the other hand, correspond to a hyperbolic partial differential equation, and could therefore be evolved in both directions.
@@ 4.3 Spatio-temporal receptive fields based on isotropic spatial smoothing
The corresponding spatio-temporal scale-space representation $L\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\times{\mathbb{R}}_{+}\times{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ of any 2+1-D video sequence $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ is then given by
@@
Differentiating the spatio-temporal kernel $T(x,t;\;s,\tau,v)$ according to (79) with respect to the spatial image coordinates $x=(x_{1},x_{2})^{T}$ and the time variable $t$ up to order 2, as well as with respect to the spatial scale parameter $s$, the temporal scale parameter $\tau$ and the elements $v_{1}$ and $v_{2}$ of the velocity vector $v$, leads to the partial derivatives shown in Figure 7 in Appendix A.
@@
Here, concerning the differential relationships over the parameters $s$ and $\tau$, the corresponding partial differential equations are parabolic, and can therefore only be evolved in the corresponding positive directions. Hence, the expressions for $\partial_{s}$ and $\partial_{\tau}$ correspond to the notions of infinitesimal generators for multi-parameter semi-groups. The differential relationship over the parameters $v_{1}$ and $v_{2}$ do, on the other hand, correspond to a hyperbolic partial differential equations, and could therefore be evolved in both directions.
@@ 4.4 Spatio-temporal receptive fields based on affine Gaussian smoothing
with the corresponding spatio-temporal scale-space representation $L\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ of any 2+1-D video sequence $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$: according to
@@
Differentiating the above spatio-temporal kernel $T(x,t;\;s,\Sigma,\tau,v)$ with respect the spatial image coordinates $x=(x_{1},x_{2})^{T}$ and the time variable $t$ up to order 2, as well as with respect to the spatial scale parameter $s$, the elements $\Sigma_{11}$, $\Sigma_{12}$ and $\Sigma_{22}$ of the spatial covariance matrix $\Sigma$, the temporal scale parameter $\tau$ and the elements $v_{1}$ and $v_{2}$ of the velocity vector $v$, then gives the partial derivatives shown in Figure 8 in Appendix B.
@@
For this joint combination of the computational structures in the two partially simplified receptive field models in Sections 4.2 and 4.3, for the differential relationships over the parameters $s$, $\Sigma_{11}$, $\Sigma_{22}$ and $\tau$, it holds that the corresponding partial differential equations are parabolic, and can therefore only be evolved in the corresponding positive directions. Hence, the expressions for $\partial_{s}$, $\partial_{\Sigma_{11}}$, $\partial_{\Sigma_{22}}$ and $\partial_{\tau}$ correspond to the notions of infinitesimal generators for multi-parameter semi-groups. The differential relationships over the parameters $\Sigma_{12}$, $v_{1}$ and $v_{2}$ do, on the other hand, correspond to hyperbolic partial differential equations, and could therefore be evolved in both directions.
@@ 5 Macroscopic cascade structures for the generalized Gaussian derivative model
To avoid explicit integration of the previous infinitesimal relationships between the receptive field responses for different values of the filter parameters, an alternative way to model such relationships is in terms of macroscopic relations. For simplicity, we again start by deriving such relations for the receptive field models that are purely based on smoothing with Gaussian kernels, thus initially excluding the non-causal spatio-temporal receptive field model based on convolution with the time-causal limit kernel (4) over the temporal domain.
@@ 5.1.1 Purely spatial receptive fields based on affine Gaussian smoothing
Applied to the purely spatial smoothing operation, corresponding to convolution with the affine Gaussian kernel, it therefore follows that the scale-space representation $L(\cdot;\;s_{2},\Sigma_{2})$ at a coarser spatial scale $(s_{2},\Sigma_{2})$ can be computed from the scale-space representation $L(\cdot;\;s_{1},\Sigma_{1})$ at any finer spatial scale $(s_{1},\Sigma_{1})$ according to
@@
Due to the linearity of the corresponding operators, this cascade smoothing property extends to the receptive field responses defined by applying spatial derivative operators ${\cal D}_{x}$ to the spatial scale-space representation of the form
@@ 5.1.2 Joint spatio-temporal receptive fields based on affine Gaussian smoothing
To express a corresponding cascade smoothing property over the joint spatio-temporal domain, let us make use of the fact that the composed effect of the smoothing operation, with a 2-D either isotropic Gaussian kernel or an anisotropic affine Gaussian kernel over the spatial domain in combination with smoothing with a 1-D non-causal temporal Gaussian kernel over the temporal domain, can equivalently be obtained by smoothing with a joint spatio-temporal Gaussian kernel over the 2+1-D joint spatio-temporal domain. Thus, when using an anisotropic affine spatial Gaussian kernel, leading to a joint spatio-temporal kernel of the form (86)
@@
this operation can equivalently be described as a 2+1-D convolution with a joint spatio-temporal kernel $g_{2+1\text{-D}}\colon{\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\times{\mathbb{S}}_{+}^{3}\rightarrow{\mathbb{R}}$ of the form
$g_{2+1\text{-D}}(p;\,m_{2+1\text{-D}},\Sigma_{2+1\text{-D}})=\\ =\frac{1}{(2\pi)^{3/2}\sqrt{\det{\Sigma_{2+1\text{-D}}}}}\,e^{-(p-m_{2+1\text{-D}})^{T}\Sigma_{2+1\text{-D}}^{-1}(p-m_{2+1\text{-D}})/2}$ (102)
@@
$\begin{split}\Sigma_{2+1\text{-D}}=\frac{\int_{p\in{\mathbb{R}}^{2+1}}p\,p^{T}\,T(p;\;s,\Sigma,\tau,v)\,\,dp}{\int_{p\in{\mathbb{R}}^{2+1}}T(p;\;s,\Sigma,\tau,v)\,dp}-m_{2+1\text{-D}}\,m_{2+1\text{-D}}^{T},\end{split}$ (105)
@@
Thus, let us assume that we have two configurations of parameter settings, that would correspond to a cascade smoothing property over the filter parameters for two joint spatio-temporal smoothing kernels of the form (101), with
  • the joint convolution kernel $T(\cdot,\cdot;\;s_{i},\Sigma_{i},\tau_{i},v_{i})$ in the lower layer with the velocity vector $v_{i}=(v_{i,1},v_{i,2})^{T}$ and the elements of the spatial covariance matrix $\Sigma_{i}$ being $\Sigma_{i,11}$, $\Sigma_{i,12}$ and $\Sigma_{i,22}$, and with
  • the joint convolution kernel $T(\cdot,\cdot;\;s_{j},\Sigma_{j},\tau_{j},v_{j})$ in the higher layer with the velocity vector $v_{j}=(v_{j,1},v_{j,2})^{T}$ and the elements of the spatial covariance matrix $\Sigma_{j}$ being $\Sigma_{j,11}$, $\Sigma_{j,12}$ and $\Sigma_{j,22}$.
@@
  • an incremental joint spatio-temporal convolution kernel $T(\cdot,\cdot;\;\Delta s,\Delta\Sigma,\Delta\tau,\Delta v)$ with the velocity vector $\Delta v=(\Delta v_{1},\Delta v_{2})^{T}$ and the elements of the spatial covariance matrix $\Delta\Sigma$ being $\Delta\Sigma_{11}$, $\Delta\Sigma_{12}$ and $\Delta\Sigma_{22}$,
@@
thus implying an explicit cascade smoothing property of the form
@@
Solving for the parameters $\Delta s$, $\Delta\Sigma_{11}$, $\Delta\Sigma_{12}$, $\Delta\Sigma_{22}$, $\Delta\tau$, $\Delta v_{1}$ and $\Delta v_{2}$ of the incremental spatio-temporal convolution kernel $T(\cdot,\cdot;\;\Delta s,\Delta\Sigma,\Delta\tau,\Delta v)$, then gives
@@ 5.1.3 Joint spatio-temporal receptive fields based on isotropic Gaussian smoothing
In relation to the results from the previous section, it is worth observing that if the spatial smoothing operation would instead only be performed based on an isotropic Gaussian kernel, with the spatial covariance matrix $\Sigma$ equal to the unit matrix $\Sigma=I$, then the parameters $(s,\tau,v)$ in the spatio-temporal receptive fields would only span 4 out of the general 6 degrees of freedom in the joint spatio-temporal covariance matrix $\Sigma_{2+1\text{-D}}$ according to (106). Thereby, the configurations for which it would be possible to implement a corresponding cascade smoothing property
@@
would on the other hand be highly degenerate. Thus, beyond the previously mentioned closedness property under anisotropic spatial affine transformations, the use of an affine Gaussian kernel for spatial smoothing also offers clear advantages in terms of the ability for an idealized vision system, to implement spatio-temporal cascade smoothing properties for the receptive fields, in order to decrease the amount of computational work in filter bank implementations of joint spatio-temporal receptive fields.